# ReX: Causal Discovery based on Machine Learning and Explainability techniques

### v1

@@ Abstract
Explainability techniques hold significant potential for enhancing the causal discovery process, which is crucial for understanding complex systems in areas like healthcare, economics, and artificial intelligence. However, no causal discovery methods currently incorporate explainability into their models to derive causal graphs. Thus, in this paper we explore this innovative approach, as it offers substantial potential and represents a promising new direction worth investigating. Specifically, we introduce ReX, a causal discovery method that leverages machine learning (ML) models coupled with explainability techniques, specifically Shapley values, to identify and interpret significant causal relationships among variables.
@@ 1 Introduction
Causal discovery —the process of identifying cause-and-effect relationships from observational data— is a pivotal challenge in artificial intelligence (AI) and machine learning. Unveiling causal structures enables robust predictions, facilitates counterfactual reasoning, and enhances decision-making processes in complex systems [1]. Traditional methods for causal discovery often rely on statistical tests for independence and structural equation modeling, which may not scale efficiently with high-dimensional data or effectively capture intricate non-linear relationships [2, 3].
In recent years, machine learning models, particularly deep learning architectures, have achieved remarkable success in predictive tasks. However, these models are typically considered “black boxes” due to their lack of interpretability. This opacity has led to a growing interest in explainable AI (XAI) techniques, with Shapley values emerging as a prominent method for interpreting model predictions [4]. Shapley values, grounded in cooperative game theory, provide a principled approach to attributing the contribution of each feature to the output of a model by quantifying the average marginal contribution of a feature across all possible subsets of features [5].
@@
Despite the advancements in both causal discovery and explainable AI, there is a notable absence of methods that integrate explainability techniques—specifically Shapley values—into the causal discovery process to extract causal graphs. This highlights a significant gap in the intersection of these two fields, offering a promising new direction.
In this paper, we propose a novel method coined ReX that integrates Shapley values into the causal discovery process, aiming to uncover causal structures by interpreting feature attributions from machine learning models. By cautiously exploiting the conceptual connection between Shapley values and causal relationships, ReX helps narrow down the set of features that play significant roles in the causal graph. This targeted approach allows us to focus on the most influential variables for further causal analysis, improving the efficiency and accuracy of causal discovery in complex datasets.
We demonstrate ReX’s competitive performance on synthetic datasets traditionally employed to evaluate new models, encompassing diverse generation processes, as well as on a complex real-world dataset. This performance underscores its practical applicability and the advantages over existing methods in addressing real-world causal discovery challenges. By integrating explainability techniques into causal discovery, we contribute to bridging the gap between predictive modeling and causal inference—a topic of increasing importance in the AI community [8, 9, 10]. This integration not only enhances the interpretability of causal models but also improves the identification of causal relationships by focusing on features with significant contributions.
@@ 1.1 Related work
Constraint-based methods, such as the PC algorithm [2] and Fast Causal Inference (FCI) [11], rely on conditional independence tests to infer causal structures. These methods explore statistical dependencies in the data to determine causal relationships between variables.
@@
Score-based methods, like Greedy Equivalence Search (GES) [12], search for the model that best fits the data according to a scoring criterion. This approach involves evaluating different graph structures and selecting the one that optimizes an objective function, balancing model fit and complexity.
@@
Structural causal models (SCMs) provide a framework for representing and estimating causal relationships through equations that describe how variables influence one another [1]. Within this category, the LiNGAM algorithm [13] exploits non-Gaussianity in data to identify causal structures in linear models, assuming linear, non-Gaussian, acyclic relationships. Causal Additive Models (CAM) [14] extend these methods by modeling nonlinear relationships using additive noise models, allowing for high-dimensional order search and penalized regression techniques. However, these models often require strong assumptions about the data-generating process and may struggle with complex nonlinearities inherent in real-world data [3].
Hybrid methods combine elements of both constraint-based and score-based approaches to improve accuracy and computational efficiency [15]. These methods aim to leverage the strengths of each approach to overcome individual limitations.
@@
Recent advancements have led to the application of machine learning techniques in causal discovery. Algorithms like NOTEARS [16] and Structural Agnostic Modeling (SAM) [17] formulate causal discovery as continuous optimization problems, enabling the use of gradient-based methods to learn directed acyclic graphs (DAGs). SAM, in particular, employs adversarial learning to model causal structures without making strong assumptions about the data distribution. These methods improve scalability and can capture nonlinear relationships but may still face challenges related to interpretability and computational demands [18].
@@ 2 Preliminaries
Causal discovery aims to infer cause-effect relationships between variables in observational data, typically represented by a causal graph. A causal graph, often modeled as a Directed Acyclic Graph (DAG), consists of nodes representing variables and directed edges indicating causal influences. The primary challenge in causal discovery is to identify these causal relationships without direct evidence from interventions, relying instead on specific assumptions about the data and its underlying generative processes.
Let $\mathbf{X}=\{X_{1},X_{2},\dots,X_{p}\}$ represent a vector of $p$ continuous variables, with unknown joint probability distribution $P(\mathbf{X})$. The observational causal discovery setting considers $m$ independent and identically distributed (iid) samples drawn from $P(\mathbf{X})$. The goal is then to infer the causal DAG $\mathcal{G}=(V,E)$, where $V$ is the set of variables and $E$ is the set of directed edges $(X_{j}\rightarrow X_{i})$ representing the causal relationship “Xj causes Xi”.
@@
$P(X_{1},X_{2},\dots,X_{n})=\prod_{i=1}^{p}P(X_{i}|\text{pa}(i,\mathcal{G})),$ (1)
@@
$X_{i}=\mkern-2.0muf\mkern-3.0mu_{i}(X_{pa(i,\mathcal{G})},\varepsilon_{i}),\ \text{with}\ \varepsilon_{i}\sim\mathcal{N}(0,1),\ \text{for}\ i=1,\dots,p,$ (2)
where $\mkern-2.0muf\mkern-3.0mu_{i}$ is a function from $\mathbb{R}^{|pa(i,\mathcal{G})|+1}\rightarrow\mathbb{R}$, and $\varepsilon_{i}$ is a unit centered Gaussian noise. SEM is a framework for modeling causal relationships where each variable is expressed as a function of its direct causes (parents) and an error term, typically assumed to be independent noise. SEMs are also essential for causal discovery because they provide a formal way to represent and quantify the effects of interventions and establish the relationships between variables.
@@ 2.1 SHAP values and causal relationships
In a causal context, if a feature does not significantly change the prediction of a target variable when added to a subset of other features, it may indicate conditional independence between the feature and the target variable, given that subset. This idea aligns with causal inference, where identifying conditional independence is key to determining whether a variable has a direct causal effect or if the relationship is mediated by other variables. By interpreting Shapley values in this way, we can gain insights into the underlying causal relationships, although Shapley values alone do not establish causality.
Several studies have explored the potential connections between SHAP values and causal relationships among variables [19, 20, 21]. These papers suggest that while SHAP values, derived from cooperative game theory, are primarily used for model explainability by attributing the contribution of each feature to a model’s prediction, they may also have implications for understanding causal relationships. However, as mentioned in Section 1, it is important to emphasize that SHAP values themselves are not inherently causal. They do not establish causal relationships but can be used in conjunction with causal inference techniques to gain insights into potential causal dynamics within the data. This area of research is ongoing, and while promising, it is crucial to distinguish between correlation, as typically highlighted by SHAP values, and causation, which requires more rigorous methodological approaches.
@@ 2.2 Mathematical Foundation
Let $F$ be the set of all features (variables), excluding the target variable $X_{i}$. Each feature $X_{j}\in F$ is considered a “player” in a cooperative game where the objective is to predict $X_{i}$. We define a function $\hat{f}(S):2^{|F|}\rightarrow\mathbb{R}$ that maps a subset of features $S\subseteq F$ to a real number representing the contribution of $S$ to predicting $X_{i}$, with $|F|$ being the cardinality of $F$. The, the Shapley value $\phi_{j}$ for feature $X_{j}$ is defined as:
@@ 2.2.1 Marginal Contributions and Conditional Independence
  • If $X_{j}$ provides no additional information about $X_{i}$ given $S$, then $X_{i}$ is conditionally independent of $X_{j}$ given $S$, denoted as $X_{i}\perp\!\!\!\perp X_{j}\mid S$. In this case $p(X_{i}\mid X_{j},S)=p(X_{i}\mid S)$, and the marginal contribution is zero ($\delta_{j,S}=0$).
[…]
@@ 2.2.2 Aggregating Over All Subsets
The Shapley value $\phi_{j}$ aggregates the marginal contributions $\delta_{j,S}$ over all subsets $S$ of $F_{\backslash\{X_{j}\}}$, weighted by the Shapley weights:
@@
2.2.3 Relating Shapley Values to Conditional Dependence
To formalize the relationship between Shapley values and conditional independence, we introduce the indicator function $I_{j,S}$:
@@ 2.2.4 Weighted Probability of Conditional Dependence
$\phi_{j}\approx\overline{\delta_{j}}\cdot P_{\text{weighted}}(X_{j}\text{ is conditionally dependent on }X_{i}).$ (10)
This approximation indicates that the Shapley value $\phi_{j}$ is proportional to both: i) the average strength of the dependence ($\overline{\delta_{j}}$); and ii) the weighted probability that $X_{j}$ is conditionally dependent on $X_{i}$.
@@ 2.2.5 Implications for Causal Discovery
By leveraging this theoretical foundation, it is correct to utilize Shapley values to estimate the conditional dependence structure among variables. However, it is important to acknowledge the assumptions underlying this connection:
  • The approximation given in Eq. (10) assumes that $\overline{\delta_{j}}$ is relatively consistent across subsets where $X_{j}$ is conditionally dependent on $X_{i}$.
  • In practice, the marginal contributions $\delta_{j,S}$ may vary significantly across different subsets $S$, and the assumption of a constant $\overline{\delta_{j}}$ serves as a simplifying approximation.
  • The predictive function $\hat{f}(S)$ is assumed to accurately capture the true conditional probabilities $p(X_{i}\mid S)$.
  • To interpret SHAP values in a causal discovery context, two core assumptions need to be assumed: faithfulness and causal sufficiency. These assumptions are essential to ensure that the inferred relationships are not merely statistical dependencies but can be interpreted as causal connections in the resulting causal DAG.

### v2

@@ Abstract
Explainable Artificial Intelligence (XAI) techniques hold significant potential for enhancing the causal discovery process, which is crucial for understanding complex systems in areas like healthcare, economics, and artificial intelligence. However, no causal discovery methods currently incorporate explainability into their models to derive the causal graphs. Thus, in this paper we explore this innovative approach, as it offers substantial potential and represents a promising new direction worth investigating. Specifically, we introduce ReX, a causal discovery method that leverages machine learning (ML) models coupled with explainability techniques, specifically Shapley values, to identify and interpret significant causal relationships among variables.
@@ 1 Introduction
Causal discovery –the process of identifying cause-and-effect relationships from observational data– is a pivotal challenge in artificial intelligence (AI) and machine learning (ML). Unveiling causal structures enables robust predictions, facilitates counterfactual reasoning, and enhances decision-making processes in complex systems [1]. Traditional methods for causal discovery often rely on statistical tests for independence and structural equation modeling, which may not scale efficiently with high-dimensional data or effectively capture intricate non-linear relationships [2, 3].
In recent years, ML models, particularly deep learning architectures, have achieved remarkable success in predictive tasks. However, these models are typically considered “black boxes” due to their lack of interpretability. This opacity has led to a growing interest in explainable AI (XAI) techniques, with Shapley values emerging as a prominent method for interpreting model predictions [4]. Shapley values, grounded in cooperative game theory, provide a principled approach to attributing the contribution of each feature to the output of a model by quantifying the average marginal contribution of a feature across all possible subsets of features [5].
Explainable Artificial Intelligence (XAI) refers to a broad set of techniques aimed at improving the transparency of ML models, commonly divided into two categories: (i) models that are interpretable by design, and (ii) post-hoc explanation methods that analyze complex, often opaque, predictive models after training. This work focuses specifically on post-hoc feature attribution. We rely on Shapley values [5] for this purpose, due to their firm theoretical foundation —particularly their satisfaction of axioms such as efficiency, symmetry, and additivity [5, 6]— as well as the availability of efficient, model-specific implementations [4, 7]. It is important to clarify that feature attributions derived from XAI methods, including Shapley values, should not be interpreted as causal explanations [8, 9].
@@
Despite the advancements in both causal discovery and explainable AI, there is a notable absence of methods that integrate explainability techniques–specifically Shapley values–into the causal discovery process to extract causal graphs. This highlights a significant gap in the intersection of these two fields, offering a promising new direction.
In this paper, we propose a novel method coined ReX that integrates Shapley values into the causal discovery process, aiming to uncover causal structures by interpreting feature attributions from ML models. By cautiously exploiting the conceptual connection between Shapley values and causal relationships, ReX helps narrow down the set of features that play significant roles in the causal graph. This targeted approach allows us to focus on the most influential variables for further causal analysis, improving the efficiency and accuracy of causal discovery in complex datasets. ReX (see Sections 2.2 and 6.1) makes use of Shapley values under the assumptions of Causal Markov Condition (Definition 1) and faithfulness (Definition 2) as a quantitative approximation of conditional dependence. This allows ReX to identify candidate causal parents from observational data, not by treating the explanations as causal per se, but by integrating XAI tools into a broader causal discovery framework.
We demonstrate ReX’s competitive performance on synthetic datasets (AppendixG) traditionally employed to evaluate new models, encompassing diverse generation processes, as well as on two real-world complex system datasets. This performance underscores its practical applicability and the advantages over existing methods in addressing real-world causal discovery challenges.
By integrating explainability techniques into the causal discovery workflow ReX aims to generate more robust causal graphs, which are fundamental for advancing from associative predictive modeling towards causal inference (i.e., elucidating cause-effect mechanisms and predicting interventional outcomes). The benefits of this specific integration within ReX are twofold: first, it enhances the transparency of the causal discovery process itself, as the selection of potential causal links is guided by interpretable Shapley values that quantify feature contributions. Second, by focusing the search on these significant and interpretable contributions, ReX can more accurately identify genuine causal relationships from observational data, leading to more reliable causal models.
@@ 1.1 Related work
Constraint-based methods, such as the PC algorithm [2] and Fast Causal Inference (FCI) [12], traditionally build causal structures by iteratively performing statistical conditional independence (CI) tests on the observational data to remove edges from an initially complete graph. These methods explore statistical dependencies in the data to determine causal relationships between variables. Their reliance on direct CI tests applied to the data for edge pruning differs from ReX’s approach. ReX first employs machine learning models to predict each variable (as detailed in Section 3.1) and then uses Shapley values (Section 2.2) derived from these models as a more holistic proxy for conditional dependence, assessing feature importance across numerous predictive contexts rather than through specific CI tests for initial graph construction.
@@
Score-based methods, like Greedy Equivalence Search (GES) [13], search for the model that best fits the data according to a scoring criterion. This approach involves evaluating different graph structures and selecting the one that optimizes an objective function, balancing model fit and complexity. This contrasts with ReX ’s methodology of constructing the graph from locally significant parent-child relationships identified via Shapley-based feature contributions (Algorithm 2), rather than optimizing a global graph score.
@@
SCMs provide a framework for representing and estimating causal relationships through equations that describe how variables influence one another [1]. Within this category, the LiNGAM algorithm [14] exploits non-Gaussianity in data to identify causal structures in linear models, assuming linear, non-Gaussian, acyclic relationships. ReX, however, uses flexible ML models (Section 3.1) such as DFNs and GBTs that inherently capture non-linearities and does not depend on non-Gaussianity for its core Shapley-based parent selection, reserving additive noise model assumptions primarily for a later edge orientation stage (Section 3.3). Causal Additive Models (CAM) [15] extend these methods by modeling nonlinear relationships using additive noise models, often utilizing techniques like penalized regression for graph estimation. While ReX also accommodates non-linearities through its underlying ML regressors and employs an additive noise assumption for orientation, its parent identification process (Algorithm 2) is driven by Shapley-based feature importance derived from these general regressors, not by fitting a specific global additive model structure with penalization for graph selection. However, these classical SCM approaches often require strong assumptions about the data-generating process and may struggle with complex nonlinearities inherent in real-world data [3].
Hybrid methods, such as Max-Min Hill-Climbing (MMHC) [16], combine elements of both constraint-based and score-based approaches, for instance, by using CI tests to restrict the search space for a subsequent score-based optimization phase. These methods aim to leverage the strengths of each approach. ReX’s integration of ML and explainability is distinct, focusing on using XAI-derived feature importances from predictive models as the primary input for parent discovery before specific graph optimization or CI testing steps.
@@
Recent advancements have led to the application of machine learning techniques in causal discovery. Algorithms like NOTEARS [17] formulate causal discovery as a continuous optimization problem by defining an algebraic acyclicity constraint, enabling gradient-based learning of the entire DAG structure simultaneously. This unified optimization differs from ReX ’s sequential workflow, which involves distinct stages for predictive modeling, Shapley-based parent discovery (Algorithm 1), edge orientation (Section 3.3), and cycle resolution (Section 3.4). Structural Agnostic Modeling (SAM) [18], in particular, employs adversarial learning to model causal structures without making strong assumptions about the data distribution. These methods improve scalability and can capture nonlinear relationships but may still face challenges related to the interpretability of the discovery process itself [19], a domain ReX specifically addresses through its inherent use of XAI techniques to guide graph construction.
@@ 2 Preliminaries
Let $\mathbf{X}=\{X_{1},X_{2},\ldots,X_{p}\}$ represent a vector of $p$ continuous variables, with unknown joint probability distribution $P(\mathbf{X})$. The observational causal discovery setting considers $m$ independent and identically distributed (i.i.d.) samples drawn from $P(\mathbf{X})$. The goal is then to infer the causal directed acyclic graph (DAG) $\mathcal{G}=(V,E)$, where $V$ is the set of variables and $E$ is the set of directed edges $(X_{j}\rightarrow X_{i})$ representing the causal relationship “Xj causes Xi”.
Causal relation vs. statistical dependence
Throughout the paper we distinguish two notions:
  1. (i)
    Statistical dependence Two variables $X$ and $Y$ are (marginally or conditionally) dependent if their joint distribution under passive observation does not factorize, e.g., $X\!\not\!\perp\!\!\!\perp Y\mid Z$. Dependence is a purely observational property.
  2. (ii)
    Causal relationship We say $X$ is a cause of $Y$ (denoted $X\!\to\!Y$ in the causal graph) when intervening on $X$ changes the distribution of $Y$: $P(Y\,|\,\operatorname{do}(X=x))$ depends on $x$ [1]. Causation implies a directed edge in the underlying interventional DAG, but does not require $X$ and $Y$ to be statistically dependent (collider structures are the standard counter-example).
@@
$P(X_{1},X_{2},\ldots,X_{n})=\prod_{i=1}^{p}P(X_{i}|\text{Pa}(i,\mathcal{G})),$ (1)
@@
Definition 1 (Causal Markov property).
$P$ is Markov with respect to $\mathcal{G}$ if each variable is independent of its non-descendants, conditioned on its parents:
$X_{i}\!\perp\!\!\!\perp\!\operatorname{nd}(X_{i})\,\big|\,\operatorname{Pa}(X_{i}),\qquad\forall\,i\in\{1,\ldots,p\}.$
@@
Definition 2 (Faithfulness).
$P$ is faithful to $G$ if every conditional independence that holds in $P$ is entailed by the DAG via $d$-separation:
$A\perp\!\!\!\perp B\,\big|\,C\text{ in }P\;\Longrightarrow\;A\;\text{and}\;B\;\text{are $d$-separated by}\;C\text{ in }G,\quad\forall\,A,B,C\subseteq V.$
Equivalently, the set of conditional independences of $P$ coincides with the set implied by $G$.
@@
$X_{i}=\mkern-2.0muf\mkern-3.0mu_{i}(X_{\text{Pa}(i,\mathcal{G})},\varepsilon_{i}),\ \text{with}\ \varepsilon_{i}\sim\mathcal{N}(0,1),\ \text{for}\ i=1,\ldots,p,$ (2)
where $\mkern-2.0muf\mkern-3.0mu_{i}$ is a function from $\mathbb{R}^{|\text{Pa}(i,\mathcal{G})|+1}\rightarrow\mathbb{R}$, and $\varepsilon_{i}$ is a unit centered Gaussian noise. SEM is a framework for modeling causal relationships where each variable is expressed as a function of its direct causes (parents) and an error term, typically assumed to be independent noise. SEMs are also essential for causal discovery because they provide a formal way to represent and quantify the effects of interventions and establish the relationships between variables.
Causal discovery aims to infer cause-effect relationships between variables in observational data, typically represented by a causal graph. A causal graph, often modeled as a DAG, consists of nodes representing variables and directed edges indicating causal influences. The primary challenge in causal discovery then, is to identify these causal relationships without direct evidence from interventions, relying instead on specific assumptions about the data and its underlying generative processes.
Definition 3 (Causal-discovery problem).
Given $m$ i.i.d. samples from an unknown $P$ that is Markov and faithful to some (unknown) DAG $\mathcal{G}$, causal discovery asks for an estimator that, as $n\to\infty$, identifies with high probability the Markov-equivalence class
$\mathcal{M}(P)\;=\;\bigl\{\,\mathcal{G}\;\big|\;P\;\text{is Markov and faithful to }\mathcal{G}\bigr\},$ (3)
or a representative DAG in that class.
Throughout the paper we work under these standard assumptions ([1, 2]); Section 4 empirically evaluates how the proposed ReX algorithm performs when the assumptions are met or mildly violated.
While causal discovery leverages patterns of statistical (in)dependence under the Markov and Faithfulness assumptions, the terms causal relation and statistical dependence are not interchangeable. The synthetic experiments in Appendix A (confounder, chain, collider) illustrate cases where high marginal dependence hides a non-causal variable and vice-versa.
@@ 2.1 SHAP values and causal relationships
When the predictor is sufficiently expressive, adding a variable $X_{j}$ to an existing feature set $S$ improves the out-of-sample prediction of the target $Y$ only if $X_{j}$ contributes information about $Y$ that is not already contained in $S$. Hence, a near-zero marginal improvement suggests the conditional independence $X_{j}\perp\!\!\!\perp Y\mid S$. This link between lack of predictive gain and conditional independence underlies many causal-discovery algorithms, which use such independencies to distinguish direct causal links from indirect (mediated) or spurious (confounded) associations when building a causal graph.
Related SHAP–causality studies
Several recent papers have examined how Shapley attributions might interface with causal reasoning [20]. introduces causal SHAP, which assumes a known DAG and re-weights coalitions accordingly; [21] applies standard SHAP values to a domain-specific regression task (coal-bed-methane wells) and discuss causal plausibility qualitatively; and [22] surveys ways in which explainers may hint at causality, but stops short of a data-driven discovery algorithm. In contrast, the present work derives an explicit link between aggregated SHAP values and conditional dependence (Section 2.2, Eq. 11) and embeds that link in a full pipeline (Sections 34) that discovers the causal graph directly from observational data without prior causal knowledge.
@@ 2.2 Mathematical Foundation
Let $F$ be the set of all features (variables), excluding the target variable $X_{i}$. Each feature $X_{j}\in F$ is considered a “player” in a cooperative game where the objective is to predict $X_{i}$. We define a function $\hat{f}(S):2^{|F|}\rightarrow\mathbb{R}$ that maps a subset of features $S\subseteq F$ to a real number representing the contribution of $S$ to predicting $X_{i}$, with $|F|$ being the cardinality of $F$. The classical definition of the Shapley value $\phi_{j}$ for feature $X_{j}$ is given by:
@@ 2.2.1 Marginal Contributions and Conditional Independence
  • If $X_{j}$ provides no additional information about $X_{i}$ given $S$, then $X_{i}$ is conditionally independent of $X_{j}$ given $S$, denoted as $X_{i}\mathbin{\perp\!\!\!\perp}X_{j}\mid S$. In this case $p(X_{i}\mid X_{j},S)=p(X_{i}\mid S)$, and the marginal contribution is zero ($\Delta_{j,S}=0$).
  • If $X_{j}$ provides additional predictive information about $X_{i}$ given $S$, then $X_{i}$ is conditionally dependent on $X_{j}$ given $S$, denoted as $X_{i}\not\!\perp\!\!\!\perp X_{j}\mid S$. In this case:
@@ 2.2.2 Aggregating over all subsets
[…]
@@
Remark.
Equations (4) and (6) are reproduced verbatim from the original work of Shapley (1953) [5]; we include them only to fix notation.
2.2.3 Shapley-weighted Conditional Dependence Indicator
To formalize the relationship between Shapley values and conditional independence, we introduce a new indicator function $I_{j,S}$:
@@ 2.2.4 Weighted probability of conditional dependence
$\phi_{j}\approx\overline{\Delta_{j}}\cdot P_{\text{weighted}}(X_{j}\text{ is conditionally dependent on }X_{i}),$ (11)
where $P_{\text{weighted}}(\cdot)$ abbreviates the Shapley weighting over coalitions. Equation (11) states that a feature’s Shapley value $\phi_{j}$ is proportional to both the average dependence strength $\Delta_{j}$ and the Shapley-weighted probability that $X_{j}$ is conditionally dependent on $Y$. [23] showed that a single Shapley summand $f(S\cup\{X_{j}\})-f(S)$ vanishes if and only if $X_{j}\perp\!\!\!\perp Y\,|\,S$ in a faithful Bayesian network. In contrast, Eqs. (6)–(10) aggregate all summands under the original Shapley weights, yielding the Shapley-weighted probability of conditional dependence in Eq. (10)—a quantity that did not appear in [23] and underpins the causal-discovery criterion of the present work.
Controlled synthetic experiments (Appendix A, Table 5) support this claim across confounder, chain, collider, and collinear structures (see Section 2.2.6).
@@ 2.2.5 Implications for causal discovery
Leveraging this theoretical foundation, Shapley values can be utilized to estimate the conditional dependence structure among variables. However, it is important to acknowledge the assumptions underlying this connection, particularly for the approximation in Eq. (11). This approximation presumes that the average marginal contribution $\overline{\Delta_{j}}$ is relatively consistent across subsets where $X_{j}$ is conditionally dependent on $X_{i}$, though in practice, individual marginal contributions $\Delta_{j,S}$ can vary significantly, making the constant $\overline{\Delta_{j}}$ a simplification. Furthermore, the predictive function $\hat{f}(S)$ is assumed to accurately capture true conditional probabilities $p(X_{i}|S)$. Finally, for a causal interpretation of the SHAP-derived insights within a discovery context, the core assumptions of faithfulness and causal sufficiency are necessary to ensure that inferred relationships represent causal connections rather than mere statistical dependencies.
@@
2.2.6 Divergence between dependence and Shapley
The theoretical connection articulated above, particularly Equation (11), is supported by controlled synthetic experiments (see Appendix A). These experiments, conducted across canonical causal structures (confounders, chains, colliders, and collinear parents), highlight several key insights:
Divergence from single-set conditional independence (CI) tests. Shapley values differ notably from traditional single-set CI tests (e.g., Fisher-Z tests). For example, in the confounder ($Z\rightarrow X$, $Z\rightarrow Y$) and chain ($X\rightarrow Z\rightarrow Y$) structures predicting $Y$, a feature $X$ may appear conditionally independent of $Y$ given $Z$ (yielding a non-significant Fisher-Z p-value, i.e., $X\mathbin{\perp\!\!\!\perp}Y\mid Z$), yet still receive substantial non-zero Shapley values. This occurs because Shapley values aggregate contributions across multiple feature coalitions, including subsets not conditioning on $Z$, thereby capturing broader predictive relevance.
Robustness under high correlation scenarios: In scenarios involving highly collinear predictors (e.g., $X\_1\approx X\_2\rightarrow Y$), Shapley values exhibit greater stability and interpretability compared to traditional CI tests or regression coefficients, which often become unstable or ambiguous. Empirical results consistently show stable Shapley values for collinear features, reflecting intuitive allocation of feature importance, consistent with properties of interventional explainers like TreeExplainer (discussed in Section 6.2).
Sensitivity to true causal drivers: Features that act as direct causes or strong mediators consistently receive high Shapley values, accurately reflecting high conditional dependence probabilities. For instance, mediators and direct causal features in confounder, chain, or collider structures consistently display elevated Shapley values, highlighting their predictive importance.
Overall, these empirical findings underscore that Shapley values, despite being model-dependent, provide comprehensive and robust assessments of feature contributions by aggregating across all feature subsets, making them valuable for causal discovery methods like ReX under assumptions of faithfulness and causal sufficiency.